Minutes of the 3rd Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Steven Watterson

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer),

Absent(with apologies) Mark Dennehy (Secretary/Match Secretary), Conor McDermottroe (Public Relations Officer),

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 27 June 2001 1800 – 2000

Distribution DURC Committee

Overview of topics

1 Minutes

1 Minutes 1 2 Meeting Time 1 3 Treasurer’s report 1 4 Chief Range Officer 2 5 Ammo & Equipment report 2 6 DURC events calender 2 7 DURC 50m Open 2 8 Junior squad coaching 2 9 Friday’s Fasseroe session 2 10 TTT 2 11 Perpetual trophy retrieval 2 12 Medal replacement 2 13 Construction of 3P air rifle frames 3 14 Time and place of next meeting 3 15 Any other Business 3

1 Minutes

The minutes from the previous meeting were approved.

2 Meeting Time

Recommendation from John aim for Mondays with meetings. General mutterings of aggreement.

3 Treasurer’s report

3 Treasurer’s report

Ben stated that we had £180 in the bank. Ben said that he’d contact Nicola for a bill for the Squad training weekend. The remaining cash will have to go on ammunition for the upcoming National championships and the Comber shoot. Henceforth, we shall pay full costs for competitions, i.e. unsubsidized entry fee and ammunition costs.

4 Chief Range Officer

New range light coming soon

5 Ammo & Equipment report

John said that he’d get some ammo. H&N Finale, T22 and Master. Due to the quantities in which we would be buying it was to be purchased from Dave Cooney. John informed us that he had a meeting with Pat Moray on Wednesday.

6 DURC events calender

Wait until NTSA calender is confirmed.

7 DURC 50m Open

Steven advocated the end of August, but we must wait until the NTSA calender is organized.

8 Junior squad coaching

John, Steven and Alan said they’d be available for coaching the potential members.

9 Friday’s Fasseroe session

John to contact Dave Cooney to secure the range for us. Steven to get bulldog clips.

10 TTT

Deal with it next week. Needs pushing at the nationals

11 Perpetual trophy retrieval

Ray’s job.

12 Medal replacement

12 Medal replacement

No money with which to replace

13 Construction of 3P air rifle frames

For standing we use the normal airgun catchers. For prone and kneeling Contact Keith. The design agreed upon was as follows:- Three pellet catchers, for three airgun firing points. The same catcher to be used for prone and kneeling. A cross section through the catcher is a right angled triangle. The vertical holds the target, perhaps made from plywood. The horizontal is similarly made from plywood and collects the pellets. The hypoteneuse is made from light steel and deflects the pellet, onto the horizontal plywood, taking the power out of it.

14 Time and place of next meeting

Monday 6pm O’Reilly Institute.

15 Any other Business

Steven to contact Jason and Deidre and find out why they are currently absent. Steven to find a copy of the curious amendment pointed out by Mark.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

15 Any other Business

Tasks agreed on at meeting